# Generative modelling for mass-mapping with fast uncertainty quantification

### v2

@@ 1 Introduction
Weak lensing has two effects to first order: convergence, $\kappa$, and shear, $\gamma$. Mass-mapping is the process of approximating the convergence from the shear and is an ill-posed inverse problem due to instrumental and noise effects. Mass-maps are incredibly useful for calculating higher-order statistics—such as Minkowski functions and bispectrum (Munshi & Coles, 2017), peak count statistics (Liu et al., 2015a; Liu et al., 2015b; Martinet et al., 2018; Harnois-Déraps et al., 2021) and scattering transform statistics (Cheng et al., 2020)—which can be compared against predictions for different cosmological models to constrain parameters and refine our understanding of the true underlying nature of the Universe.
Given the advances in observational technology, we are now in the era of precision cosmology, where weak lensing measurements can provide unprecedented insights into the large-scale structure of the Universe (Mandelbaum, 2018). Mass-mapping techniques must advance to accommodate the increasing precision of these measurements. The seminal, and still most widely-used mass-mapping method is the Kaiser-Squires technique (Kaiser & Squires, 1993), which directly inverts the noisy shear field. Although popular for its speed and computational efficiency, this method does not account for instrumental effects and noise, resulting in reconstructions that are noisy to the point of requiring mandatory post-processing. This post-processing typically involves a Gaussian smoothing of the convergence map, which leads to a loss of small-scale structure in the map. The Wiener filter (Wiener, 1949), a more sophisticated linear method, assumes a Gaussian prior on the convergence and serves as the maximum-a-posteriori (MAP) estimator of the problem, though it typically requires selecting a fiducial cosmology. Wavelet-based methods (Lanusse et al., 2016; Price et al., 2019; Price et al., 2020; Price et al., 2021a; Price et al., 2021b; Starck et al., 2021) aim to retain small-scale structure through the use of wavelet-based priors. These priors have led to increased performance, however, the lack of flexibility with hand-crafted priors has led to growing interest in deep learning methods, which can learn the priors from data itself.
@@
Recent deep learning approaches for mass-mapping fall into two main categories. The first involves using deep learning for post-processing. For example Jeffrey et al. (2020) introduce a technique to post-process a reconstructed convergence field using a convolutional U-Net (Ronneberger et al., 2015). This approach is very fast but does not provide any uncertainty quantification. Conversely, Shirasaki et al. (2021) use conditional generative adversarial networks (GANs) to learn noise maps which can then be used to denoise convergence maps. Again, uncertainties are not provided. The second general category seeks to directly learn the posterior distribution of the data using deep learning. For example, Remy et al. (2023) introduce a technique that generates samples from a learned approximate posterior distribution, using data-driven priors learned by neural score estimation. These approximate posterior samples can then be used to construct a point-estimate reconstruction and to estimate uncertainties. However, it comes at the cost of being very slow, requiring 10 GPU minutes to generate a single independent approximate posterior sample. For both general categories generative modelling techniques in particular have proven highly effective due to the rich data-driven prior information that they are able to capture (Shirasaki et al., 2021; Remy et al., 2023). While all of these deep learning methods have provided promising results, particularly in recovering both large and small-scale structure in convergence maps, there is still work to be done to develop methods that provide high fidelity reconstructions, are computationally efficient, and also provide uncertainty estimates. Fast generation of approximate posterior samples is necessary for integration into downstream cosmological parameter estimation and model comparison pipelines so that uncertainties in the mass-mapping process are captured.
@@ 2.1 Weak Gravitational Lensing
[…]
@@
[…]
@@
where $\eth$ denotes the spin-$s$ raising operator,
ð$\eth=-\sin^{s}\vartheta\big{(}\partial_{\vartheta}+\frac{i\partial_{\varphi}}{\sin\theta}\big{)}\sin^{-s}\vartheta\approx-(\partial_{x}+i\partial_{y})$ (6)
@@
ð¯$\bar{\eth}=-\sin^{-s}\vartheta\big{(}\partial_{\vartheta}-\frac{i\partial_{\varphi}}{\sin\theta}\big{)}\sin^{s}\vartheta\approx-(\partial_{x}-i\partial_{y}).$ (7)
@@
Substituting the flat-sky approximation of the $\eth$ and $\bar{\eth}$ into the expression for the shear and convergence one finds that
@@ 2.2 Lensing Inverse Problem
With observations of $\gamma$ to hand one may attempt to infer $\kappa$ by exploiting their Fourier space relationship. The most naïve algorithm by which $\kappa$ may be recovered given observations of $\gamma$ is by simply inverting this relation $\tilde{\kappa}_{\text{KS}}=\mathbf{D}^{-1}\tilde{\gamma}$, which is the original method developed by Kaiser & Squires (1993). In the absence of observation noise and masking effects $\kappa_{\text{KS}}$ is equivalent to the maximum-likelihood estimator $\kappa_{\text{MLE}}$. However, in realistic scenarios noise contributions are overwhelmingly dominant and complex masking is present, hence the two estimators are by no means equivalent.
The Kaiser-Squires estimator is known for its computational efficiency and simplicity, however, it comes with several major drawbacks. First and foremost, it does not account for the observational noise, which consequently propagates directly to the reconstruction, thereby making post-processing virtually mandatory to facilitate downstream analysis. This post-processing typically takes the form of Gaussian smoothing, which leads to the loss of non-Gaussian features in the convergence map. In particular, this results in a suppression of peaks in the reconstruction, and loss of small-scale structure—both of which are critical information for contemporary cosmology. Second, it does not provide any measure of the uncertainties associated with the reconstruction.
Since the Kaiser-Squires method was proposed there have been many other methods developed for mass-mapping, such as sparsity-based wavelet methods (Lanusse et al., 2016; Price et al., 2019; Price et al., 2020, 2021b; Starck et al., 2021) and deep learning architectures (Jeffrey et al., 2020; Shirasaki et al., 2021; Remy et al., 2023). Several of these methods have further been extended from the flat-sky to the sphere for wide-field mass-mapping (Wallis et al., 2022; Chang et al., 2018; Price et al., 2021a). Nevertheless, it is fair to say mass-mapping is by no means a solved problem.
@@ 3.3 Regularised conditional GANs
Regularised conditional GANs (Bendel et al., 2024) are a recent development, designed to overcome mode collapse in conditional GANs—which as previously mentioned is a more acute problem than with traditional GANs, and also harder to solve due to the one-to-one ${x,y}$ pairing of data. Within this framework, the generator aims to solve the following minimisation problem
@@ 4 Simulations, Training, and Validation
This section details the simulations and mock dataset used to train our model. We first discuss the $\kappa$TNG simulations, a collection of convergence maps based on the IllustrisTNG simulations, before moving on to describe how we used this weak lensing map suite to build a mock catalog of 10,000 convergence maps in the style of the COSMOS survey. This catalog was then used to train, validate, and test our model.
@@ 4.1 KappaTNG Simulations
The $\kappa$TNG simulations are a suite of $10,000$ mock weak lensing maps (Osato et al., 2021), based on the IllustrisTNG hydrodynamical simulations (Springel et al., 2018). All simulations assume the flat $\Lambda$CDM cosmology as in Planck 2015 (Ade et al., 2016), with $H_{0}=67.74\>\text{kms}^{-1}\text{Mpc}^{-1}$, baryonic density $\Omega_{b}=0.0486$, matter density $\Omega_{m}=0.3089$, and spectral index of scalar perturbations $n_{s}=0.9667$.
@@ 4.3 Mock COSMOS Dataset
In order to create mock COSMOS maps we utilised both the $\kappa$TNG simulations, and the Schrabback et al. (2010) shape catalog. As mentioned, this shape catalog is divided into a bright and faint catalog, which we combined into a full catalog. We discarded galaxies with photometric redshifts $z_{\text{phot}}<0.6$ for reasons specified in Section 4.2. Then we calculated the redshift distribution, $p(z)$, of the galaxies in the full catalog.
@@
The $\kappa$TNG maps are sliced at discrete redshifts between $z\in[0,2.568]$, leading to $40$ evenly-spaced source planes. For our mock COSMOS maps we required a redshift up to $z=5$, therefore we chose redshift values $z_{i}$ with spacing equal to the $\kappa$TNG slices. This resulted in $80$ redshift values $z_{i}$, leading to $79$ redshift bins, centered on the redshift values of the $\kappa$TNG maps. Note that the bin size is halved for the first and last bins. For $z>2.568$ we follow Remy et al. (2023) and reused the highest redshift slice convergence map ($z=2.568$), while calculating a new weight for each bin.
@@
In summary, we created $10,000$ convergence maps of size $300\times 300$ pixels. In order to convert these to mock shear maps, we used the forward model described by Equation (10). We added spatially varying noise to the shear, treating the real and imaginary components separately by calculating the standard deviation of the $\gamma_{1}$ and $\gamma_{2}$ estimates in the COSMOS shape catalog respectively. We then simulated the noise by multiplying the two standard deviations by a random normal distribution, and adding to the real and imaginary components of the clean mock COSMOS shear maps.
@@ 4.5 Model Validation
where $\text{MAX}_{I}$ is the maximum possible pixel value, and MSE is the mean squared error between the truth and the reconstruction. We calculated the PSNR of a single posterior sample as well as the PSNR of a reconstruction made from the average of $N=32$ samples (for discussion on our choice of $N$ see Section 5.1.1).
@@ 5 Results
To assess reconstruction quality, we compared to the pixel-wise absolute error $=|\hat{x}_{(N)}-x|$ between the reconstruction and the ground truth. We also used the absolute error to asses the usefulness of using the standard deviation of the generated approximate posterior samples as a measure of uncertainty.
@@ 5.1.1 Reconstructions of Simulations
Table 1 further explores how the reconstruction changes as the number of generated samples used changes. It shows the PSNR and structural similarity index measure (SSIM) (Wang et al., 2004) on reconstructions with different values of $N$, calculated during model validation. These metrics were calculated by comparing reconstructions with the ground truths for the mock data.
@@
Figure 5: Demonstration of how the $N$-sample reconstruction varies for $N\in\{1,4,8,32\}$, $N=1$ being a single posterior sample, for the zoomed-in region shown in the red box. The figure also shows the Kaiser-Squires map for the same region. As can be seen, the reconstruction becomes smoother as $N$ increases, however, the prominent features remain. An individual sample has a far higher level of detail, comparable with the true map, however, it can be seen that features differ slightly to the truth, indicating why it is necessary to average over a number of samples. Despite some loss of the smallest-scale structure for $N=32$, there is less peak suppression than the Kaiser-Squires reconstruction.
@@
Table 2: Results of validation metrics. The Pearson correlation coefficient, RMSE, and PSNR, were calculated for the Kaiser-Squires reconstruction (with $\sigma=1\,\text{arcmin}$ smoothing, chosen to minimise RMSE) and our 32-sample MMGAN reconstruction across a validation dataset. The results were averaged and then used to create this Table. Metrics for methods marked with an asterisk (*) are sourced from Table 1 in Remy et al. (2023) and therefore should not be directly compared with our results, since they consider a different validation set. Instead, they serve to provide a general comparison between MMGAN and other methods. Notably, the Kaiser-Squires results differ slightly from those reported in Remy et al. (2023), likely due to variations in the validation set.
@@
MMGAN (Ours) 0.727 0.0197 31.674
Kaiser-Squires 0.622 0.0229 30.387
@@ 5.1.2 Uncertainty Quantification Validation
A well as evaluating reconstruction quality, we also assessed the effectiveness of our uncertainty quantification. When building reconstructions of simulated convergence maps, we qualitatively compared the resulting uncertainty map with the pixel-wise absolute error between the MMGAN reconstruction and the ground truth. Visually, there is a correlation between these fields, as can be seen in Figure 3 and Figure 8.
We also calculated the empirical coverage probability of reconstructions compared to the ground truth maps at 90% and 95% confidence intervals. We found that across 100 maps from our test set, the coverage was in reasonably close agreement at 85% and 89% respectively.
@@ 6 Conclusions
Given a noisy shear observation, we construct a pseudo-reconstruction, and pass both into the MMGAN generator, which then outputs an approximate posterior sample. We take the posterior mean of $N=32$ approximate posterior samples as our final reconstruction, and the standard deviation of these samples to quantify our uncertainties. Currently, we choose a Kaiser-Squires map as our pseudo-reconstruction, however, this could be replaced with a more sophisticated reconstruction method, such as the Wiener filter. MMGAN does not require an explicit choice of cosmology, which is another reason why we chose the Kaiser-Squires methods as our pseudo-reconstruction. We trained MMGAN on a mock dataset with a fixed cosmology, although in future it would be interesting to explore how MMGAN performs when trained on a dataset with a range of cosmologies.
To train MMGAN, we used mock COSMOS-style shear and convergence maps, made from the $\kappa$TNG simulations. We validated our model on a subset of the mock COSMOS data not seen during training. We used the PSNR to ensure the model was suitably constrained in terms of variance in its sample generation, and then used standard metrics such as the PSNR, RMSE, and Pearson correlation coefficient to select the best training epoch of our model. We found that our MMGAN reconstructions are able to capture both large- and small-scale structure, and do not require any post-processing such as smoothing, which is known to suppress peaks. The resulting MMGAN model leverages data-driven priors to produce high-fidelity reconstructions with uncertainty estimates, all generated within seconds.

### v4

@@ 1 Introduction
Weak lensing has two effects to first order: convergence, $\kappa$, and shear, $\gamma$. Mass-mapping is the process of approximating the convergence from the shear and is an ill-posed inverse problem due to instrumental and noise effects. Mass-maps are incredibly useful for calculating higher-order statistics—such as Minkowski functions and bispectrum (Munshi & Coles, 2017), peak count statistics (Liu et al., 2015a; Liu et al., 2015b; Martinet et al., 2018; Harnois-Déraps et al., 2021) and scattering transform statistics (Cheng et al., 2020)—which can be compared against predictions for different cosmological models to constrain parameters and refine our understanding of the true underlying nature of the Universe. Given the advances in observational technology, we are now in the era of precision cosmology, where weak lensing measurements can provide unprecedented insights into the large-scale structure of the Universe (Mandelbaum, 2018). To fully exploit the potential of these applications, mass-mapping methods must advance to provide precise and accurate reconstructions on par with the quality of the data now available to us. These methods should aim to retain small-scale structure and minimise the loss of information due to noise. Furthermore, they should provide well-characterised uncertainties, and be computationally feasible for large-scale inference to enable robust cosmological parameter estimation and model comparison.
The Kaiser-Squires technique (Kaiser & Squires, 1993) is the seminal mass-mapping approach, based on direct inversion of the noisy shear field. It remains widely-used due to its speed and computationally efficiency. In an idealised setting it is equivalent to the maximum-likelihood estimator (MLE) due to the straightforward invertibility of the forward model. In practice, however, additional factors such as masking and instrumental effects break this invertibility, making mass-mapping an ill-posed problem (Price et al., 2021a). Prior information is often injected to regularise ill-posed problems to recover effective solutions—an aspect not addressed by the Kaiser-Squires method. Instead, it is common to include a post-processing step in the Kaiser-Squires approach, typically Gaussian smoothing, to improve the signal-to-noise ratio of a reconstruction. However, this smoothing suppresses small-scale structures, potentially discarding valuable non-Gaussian information. While this may be an acceptable trade-off in certain applications (e.g. Kovács et al. 2022), it is fundamentally limiting for higher-order statistical analyses and cosmological parameter estimation, where the small-scale structure is crucial. Wiener filtering (Wiener, 1949) provides a maximum a posteriori (MAP) estimate, assuming a Gaussian prior based on a fiducial cosmology for the convergence. Like Gaussian-smoothed Kaiser-Squires maps, the Wiener filter also suppresses small-scale structures that are dominated by noise, resulting in a similar loss of information. Wavelet-based methods (Lanusse et al., 2016; Price et al., 2019; Price et al., 2020; Price et al., 2021b; Price et al., 2021c; Starck et al., 2021) aim to retain small-scale structure through the use of wavelet-based priors. These priors have led to increased performance, however, the lack of flexibility with hand-crafted priors has led to growing interest in deep learning methods, which can learn the priors from data itself.
@@
Recent deep learning approaches for mass-mapping fall into two main categories. The first involves using deep learning for post-processing. For example, Jeffrey et al. (2020) introduce a technique to post-process a reconstructed convergence field using a convolutional U-Net (Ronneberger et al., 2015). This approach is very fast but does not provide any uncertainty quantification. Conversely, Shirasaki et al. (2021) use conditional generative adversarial networks (GANs) to learn noise maps which can then be used to denoise convergence maps. Again, uncertainties are not provided. The second general category seeks to directly learn the posterior distribution of the data using deep learning. For example, Remy et al. (2023) introduce a technique that generates samples from a learned approximate posterior distribution, using data-driven priors learned by neural score estimation. These approximate posterior samples can then be used to construct a point-estimate reconstruction and to estimate uncertainties. However, it comes at the cost of being very slow, requiring 10 GPU minutes to generate a single independent approximate posterior sample. For both categories generative modelling techniques have proven highly effective due to the rich data-driven prior information that they are able to capture (Shirasaki et al., 2021; Remy et al., 2023). While all of these deep learning methods have provided promising results, particularly in recovering both large and small-scale structure in convergence maps, there is still work to be done to develop methods that provide high fidelity reconstructions, are computationally efficient, and also provide uncertainty estimates. Fast generation of approximate posterior samples is necessary for integration into downstream cosmological parameter estimation and model comparison pipelines so that uncertainties in the mass-mapping process are captured.
@@ 2.1 Weak Gravitational Lensing
$\phi(r,\omega)=\dfrac{2}{c^{2}}\int_{0}^{r}dr^{\prime}\dfrac{f_{K}(r-r^{\prime})}{f_{K}(r)f_{K}(r^{\prime})}\Phi(r^{\prime},\omega),$ (3)
@@
$\nabla^{2}\phi(r,\omega)=\dfrac{3\Omega_{M}H_{0}^{2}}{c^{2}}\int_{0}^{r}dr^{\prime}\dfrac{f_{K}(r-r^{\prime})}{f_{K}(r)f_{K}(r^{\prime})}\dfrac{\delta(r,\omega)}{a(r)}.$ (4)
@@
[…]
$\eth=-\sin^{s}\vartheta\big(\partial_{\vartheta}+\frac{i\partial_{\varphi}}{\sin\theta}\big)\sin^{-s}\vartheta\approx-(\partial_{x}+i\partial_{y})$ (6)
@@
$\bar{\eth}=-\sin^{-s}\vartheta\big(\partial_{\vartheta}-\frac{i\partial_{\varphi}}{\sin\theta}\big)\sin^{s}\vartheta\approx-(\partial_{x}-i\partial_{y}).$ (7)
@@
[…]
@@ 2.2 Lensing Inverse Problem
With observations of $\gamma$ to hand one may attempt to infer $\kappa$ by exploiting their Fourier space relationship. The most naïve algorithm by which $\kappa$ may be recovered given observations of $\gamma$ is by simply inverting this relation $\tilde{\kappa}_{\text{KS}}=\mathbf{D}^{-1}\tilde{\gamma}$, which is the original method developed by Kaiser & Squires (1993). As discussed in the introduction, in an idealised setting $\kappa_{\text{KS}}$ is equivalent to the maximum-likelihood estimator (Price et al., 2021a). However, in realistic scenarios noise contributions are overwhelmingly dominant and complex masking is present, thus the two estimators are by no means equivalent.
The Kaiser-Squires estimator is known for its computational efficiency and simplicity, however, it comes with several major drawbacks. First and foremost, it does not account for observational noise, which consequently propagates directly to the reconstruction. In order to use the mass-maps for cosmological inference, post-processing methods are conventionally used to improve their signal-to-noise ratio. This post-processing typically takes the form of Gaussian smoothing, which leads to the loss of non-Gaussian features in the convergence map. In particular, this results in a suppression of peaks in the reconstruction, and loss of small-scale structure—both of which are critical information for contemporary cosmology. Second, it does not provide any measure of the uncertainties associated with the reconstruction.
Since the Kaiser-Squires method was proposed there have been many other methods developed for mass-mapping, such as sparsity-based wavelet methods (Lanusse et al., 2016; Price et al., 2019; Price et al., 2020, 2021c; Starck et al., 2021) and deep learning architectures (Jeffrey et al., 2020; Shirasaki et al., 2021; Remy et al., 2023). Several of these methods have further been extended from the flat-sky to the sphere for wide-field mass-mapping (Wallis et al., 2022; Chang et al., 2018; Price et al., 2021b). Nevertheless, it is fair to say mass-mapping is by no means a solved problem.
@@ 3.3 Regularised conditional GANs
Regularised conditional GANs (Bendel et al., 2024) are a recent development, designed to overcome mode collapse in conditional GANs—which as previously mentioned is a more acute problem than with traditional GANs, and also harder to solve due to the one-to-one x,y{x,y} pairing of data. Within this framework, the generator aims to solve the following minimisation problem
@@ 4 Simulations, Training, and Validation
[…]
@@ 4.1 KappaTNG Simulations
The maps were generated by creating light cones with an opening angle of $5\times 5\>\text{deg}^{2}$, from the IllustrisTNG simulations, made by stacking TNG snapshots along the line of sight. The mock weak lensing maps were then created by tracing the light cones from $z=0$ up to the target redshift, $z_{s}\in[0.00,2.57]$. To create the full suite, a large number of random flips, rotations, and translations were applied to the IllustrisTNG snapshots. The subsequent maps were shown to be statistically independent (Osato et al., 2021). Each map is of size $1024\times 1024$ pixels, with a resolution of $0.29\>\text{arcmin/pixel}$.
@@ 4.3 Mock COSMOS Dataset
[…]
@@
[…]
@@
In summary, we created $10,000$ convergence maps of size $300\times 300$ pixels. In order to convert these to mock shear maps, we used the forward model described by Equation (10). We added spatially varying noise to the shear, treating the real and imaginary components separately by calculating the standard deviation of the $\gamma_{1}$ and $\gamma_{2}$ estimates in the COSMOS shape catalog respectively. We then simulated the noise by multiplying the two standard deviations by a random normal distribution, and adding to the real and imaginary components of the clean mock COSMOS shear maps. We note that these maps do not account for instrumental effects such as photometric uncertainty, intrinsic galaxy alignment, or baryonic feedback effects. However, if these effects were incorporated into the training data, the model could, in principle, learn to account for them.
@@ 4.5 Model Validation
where $\text{MAX}_{I}$ is the maximum possible pixel value (which we set to 1), and MSE is the mean squared error between the truth and the reconstruction. We calculated the PSNR of a single posterior sample as well as the PSNR of a reconstruction made from the average of $N=32$ samples (for discussion on our choice of $N$ see Section 5.1.1).
@@ 5 Results
To assess reconstruction quality, we compared to the pixel-wise absolute error $|\hat{x}_{(N)}-x|$ between the reconstruction and the ground truth. We also used the absolute error to asses the usefulness of using the standard deviation of the generated approximate posterior samples as a measure of uncertainty.
@@ 5.1.1 Reconstructions of Simulations
Table 1 further explores how the reconstruction changes as the number of generated samples used changes. It shows the PSNR and structural similarity index measure (SSIM) (Wang et al., 2004) on reconstructions with different values of $N$, calculated during model validation. The SSIM is computed as
$\text{SSIM}(x,y)=\dfrac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})},$ (31)
where $\mu_{x}$ and $\mu_{y}$ are the means of $x$ and $y$, $\sigma_{x}$ and $\sigma_{y}$ are the standard deviations of $x$ and $y$, $\sigma_{xy}$ is the covariance of $x$ and $y$, and $C_{1}$ and $C_{2}$ are constants. These metrics were calculated by comparing reconstructions with the ground truths for the mock data.
@@
Figure 5: Demonstration of how the $N$-sample reconstruction varies for $N\in\{1,4,8,32\}$, $N=1$ being a single posterior sample, for the zoomed-in region shown in the red box. The figure also shows the Kaiser-Squires map for the same region. As can be seen, the reconstruction becomes smoother as NN increases, however, the prominent features remain. An individual sample has a far higher level of detail, comparable with the true map, however, it can be seen that features differ slightly to the truth, indicating why it is necessary to average over a number of samples. Despite some loss of the smallest-scale structure for $N=32$, there is less peak suppression than the Kaiser-Squires reconstruction.
@@
Table 2: Results of validation metrics. The Pearson correlation coefficient, RMSE, and PSNR, were calculated for the Kaiser-Squires reconstruction (with $\sigma=1\,\text{arcmin}$ smoothing, chosen to minimise RMSE) and our 32-sample MMGAN reconstruction across a validation dataset. The results were averaged and then used to create this Table. Metrics for methods marked with an asterisk (*) are sourced from Table 1 in Remy et al. (2023) and therefore should not be directly compared with our results, since they consider a different validation set. Instead, they serve to provide a general comparison between MMGAN and other methods. Notably, the Kaiser-Squires results differ slightly from those reported in Remy et al. (2023), likely due to differences in the randomly selected validation sets.
@@
MMGAN (Ours) 0.727 0.0197 34.106
Kaiser-Squires 0.619 0.0229 32.803
@@ 5.1.2 Uncertainty Quantification Validation
As well as evaluating reconstruction quality, we also assessed the effectiveness of our uncertainty quantification. When building reconstructions of simulated convergence maps, we qualitatively compared the resulting uncertainty map with the pixel-wise absolute error between the MMGAN reconstruction and the ground truth. Visually, there is a correlation between these fields, as can be seen in Figure 3 and Figure 8.
We also investigated MMGANs uncertainty maps using coverage tests, which assess whether the predicted uncertainties accurately reflect the true reconstruction error. Preliminary results suggest that MMGAN achieves strong performance. A comprehensive analysis of these tests will be presented in Whitney et al. (prep).
@@ 6 Conclusions
Given a noisy shear observation, we construct a pseudo-reconstruction, and pass both into the MMGAN generator, which then outputs an approximate posterior sample. We take the posterior mean of $N=32$ approximate posterior samples as our final reconstruction, and the standard deviation of these samples to quantify our uncertainties. Currently, we choose a Kaiser-Squires map as our pseudo-reconstruction, however, this could be replaced with a more sophisticated reconstruction method, such as the Wiener filter. MMGAN does not require an explicit choice of cosmology, which is another reason why we chose the Kaiser-Squires method as our pseudo-reconstruction. We trained MMGAN on a mock dataset with a fixed cosmology, which we acknowledge may bias the learned model towards that cosmology. However, this is a limitation of the dataset rather than the method itself—there is no fundamental reason MMGAN could not be trained on a dataset with a range of cosmologies. Additionally, we did not account for astrophysical and observational systematics such as intrinsic alignments, photometric uncertainties, and shear multiplicative biases within our training data. These effects can introduce biases in the shear measurements, potentially impacting both reconstruction accuracy and uncertainty estimated. In future work, it would be interesting to explore how MMGAN performs when trained on data that accounts for these effects, as understanding their impact on the reconstruction and uncertainty maps is an important step towards ensuring the robustness of the method.
[…]